# AlignRAG: Leveraging Critique Learning for Evidence-Sensitive Retrieval-Augmented Reasoning

### v3

@@ Abstract
Retrieval-augmented generation (RAG) has become a widely adopted paradigm for enabling knowledge-grounded large language models (LLMs). However, standard RAG pipelines often fail to ensure that model reasoning remains consistent with the evidence retrieved, leading to factual inconsistencies or unsupported conclusions. In this work, we reinterpret RAG as Retrieval-Augmented Reasoning and identify a central but underexplored problem: Reasoning Misalignment—the divergence between an LLM’s internal reasoning trajectory and the evidential constraints provided by retrieval. To address this issue, we propose AlignRAG, a novel iterative framework grounded in Critique-Driven Alignment (CDA). At the heart of AlignRAG lies a contrastive critique synthesis mechanism that generates retrieval-sensitive critiques while mitigating self-bias. This mechanism trains a dedicated retrieval-augmented Critic Language Model (CLM) using labeled critiques that distinguish between evidence-aligned and misaligned reasoning. Alignment signals for supervision are obtained through self-supervised or externally guided labeling strategies. The resulting CLM is explicitly optimized for evidence sensitivity, enabling it to detect and revise reasoning errors during inference without relying solely on self-generated feedback. Empirical evaluations show that our 8B-parameter CLM improves performance over the Self-Refine baseline by 12.1% on out-of-domain tasks and outperforms a standard 72B-parameter CLM by 2.2%, while remaining compatible with existing RAG architectures as a plug-and-play module. Moreover, AlignRAG demonstrates strong robustness under both informative and noisy retrieval scenarios, validating its ability to thrive even when retrieval quality degrades. Our source code is provided at link.
@@ 1 Introduction
Large Language Models (LLMs) have significantly advanced natural language understanding and generation capabilities. Retrieval-Augmented Generation (RAG) [1, 2, 3, 4, 5, 6, 7, 8] has emerged as a prominent paradigm for grounding LLM responses with external knowledge. However, RAG systems exhibit notable fragility, particularly when confronted with irrelevant or noisy retrieved evidence [9, 10]. Existing methods primarily rely on static, training-time optimizations, which are often insufficient to address the dynamic challenges of error propagation during inference [11, 12]. We identify a critical, yet understudied, failure mode in RAG: reasoning misalignment—a disconnect between the model’s reasoning process and the retrieved evidence. Prior work has focused predominantly on improving retrieval quality or generating more robust outputs, largely overlooking the explicit alignment of the reasoning steps with the provided evidence [13, 6, 14, 15, 16]. While reflective approaches like Self-RAG [17] attempt error detection, they often necessitate architectural modifications or task-specific fine-tuning, limiting their generalizability.
@@
In summary, this paper makes the following key contributions: (1) We reconceptualize RAG as Retrieval-Augmented Reasoning and identify Reasoning Misalignment as a fundamental, understudied failure mode. (2) We introduce critique learning for RAG, a novel pipeline for training CLMs to generate retrieval-augmented critiques while mitigating self-preference bias through a contrastive synthesis approach. We show that an 8B-parameter CLM trained with our method outperforms a vanilla 72B-parameter CLM. (3) We propose AlignRAG, the first test-time framework that utilizes CDA steps to iteratively optimize the RAG reasoning process, explicitly aligning it with evidential constraints from retrieved content. (4) We provide extensive empirical validation demonstrating that AlignRAG achieves SOTA performance and significantly improves reasoning quality and robustness across various benchmarks and retrieval scenarios.
@@
2 Related Work
RAG [25, 26, 27, 28, 29, 30, 31] enhances LLMs by grounding generation in external knowledge. While significant efforts have focused on improving retrieval accuracy [32, 33, 34, 35, 36] or training better generators [37, 38], a persistent challenge is the presence of noisy or irrelevant retrieved content. To mitigate this, recent methods filter or denoise context, predominantly via training-time optimizations [13, 39, 6, 40, 41], such as InstructRAG’s [11] self-supervised denoising. However, these approaches are limited by their static nature and fail to address dynamic error propagation during inference [12].
@@ 3 Reasoning Misalignment in RAG
Phase 3: Evidence-Integrated Synthesis. Misalignment here involves generating reasoning steps or conclusions $\text{step}_{i}\in\text{y}$ that are not logically supported by the relevant evidence $E_{i}\subseteq\mathcal{D}$ used for that step, violating the entailment $E_{i}\models\text{step}_{i}$, or creating inconsistencies when synthesizing information from multiple parts of $\mathcal{D}$ into a coherent justification [51, 12].
@@ 4.2.1 Training Corpus Construction
To model the ambiguity and diversity inherent in real-world retrieval scenarios [52, 53], we construct a structured training dataset $\mathcal{S}=\{(q_{i},a_{i},\mathcal{D}_{i},\mathbf{c}_{i})\}_{i=1}^{N}$, where each instance includes a query $q_{i}$, its gold answer $a_{i}$, a retrieved document set $\mathcal{D}_{i}$, and a context granularity vector $\mathbf{c}_{i}\in\{0,1\}^{3}$. The vector encodes three orthogonal axes of contextual variation:
@@ 4.3 Critique-Driven Alignment Systems
[…]
@@ 5.1 Experiment Setup
Dataset. To train a strong critique generator, we construct a 10K dataset by sampling 2K instances from each of five benchmarks: PopQA [59], TriviaQA [60], NaturalQuestions [61], 2WikiMultihopQA [62], and ASQA [63].
Furthermore, we evaluate our method on the same five in-domain benchmarks, along with two out-of-distribution (OOD) tasks, i.e., HotpotQA [64] and SQuAD [65].
Baselines. In our experiments, we compare our method against a range of non-retrieval and retrieval-based baselines. For non-retrieval baselines, we include Chain-of-Thought (CoT) prompting [66] applied to models without retrieval augmentation. For standard RAG, we report performance from Vanilla Reasoning [11], which performs step-by-step answer generation based on the retrieved passages. To assess the benefits of intermediate supervision, we include training-time refinement baselines such as RetRobust[52] and InstructRAG[11]. Our main comparison is with these test-time refinement methods, as they share similar objectives. For test-time refinement, we evaluate Self-RAG [17] and Self-Refine, which iteratively revises outputs based on self-generated critique.
@@ 5.3 Analysis
Figure 4: AlignRAG performance under Noisy Retrieval (4(a)) and Informative Retrieval (4(b)) scenarios.
@@
This dual strategy, as demonstrated in Figures 4(a) and 4(b), empowers robust reasoning even under these adversarial conditions. Consequently, AlignRAG (Figure 4(a)) significantly outperforms conventional methods like Vanilla RAG and Self-Refine, which falter in such scenarios, by effectively transforming imperfect external retrievals and inherent model understanding into actionable insights.
@@
of using a high-capacity Strong LLM, i.e., Qwen2.5-72B-Instruct, in our Contrastive Critique Synthesis (CCS) framework. Specifically, we compare three CLM supervision strategies under three backbones: (1) Vanilla Critique Synthesis, where the CLM is trained with critiques directly generated by Qwen2.5-72B-Instruct, (2) directly using the 72B model as the CLM, and (3) our CCS-based CLM trained with critiques synthesized by the 72B Strong LLM. As shown in Figure 5, our method consistently outperforms vanilla critique synthesis, validating the benefit of introducing contrastive reasoning signals. Remarkably, our method even surpasses the directly supervised 72B CLM in all backbones, suggesting that contrastive critique training enhances generalization and reduces reliance on model scale. These results demonstrate the strength of our approach in leveraging powerful LLMs for scalable and efficient CLM training.
@@ 5.4 Ablation Study
In this section, we conduct an ablation study across seven benchmarks to evaluate the impact of the Critic Language Model (CLM), a key component of our framework. We compare four configurations: (1) Vanilla RAG, which performs generation without critique; (2) RAG with Frozen CLM, which uses a CLM based on LLaMA3.1-8B without training; (3) RAG with Trained CLM, where the CLM is trained using critiques generated by LLaMA3.1-8B over first responses produced only by a Weak LLM, without using contrastive signal. (4) RAG with Trained CLM using our CCS, where the CLM is based on LLaMA3.1-8B and further optimized via our contrastive critique synthesis (CCS) strategy.
@@
As shown in Figure 6, incorporating even an untrained CLM leads to consistent gains over the vanilla RAG baseline, indicating that an auxiliary critic can already provide useful feedback during inference. Further improvements are achieved by critically fine-tuning the CLM with only the weak LLM used in our method, demonstrating that training the CLM on critiques over weak LLM responses enables it to learn generalizable error patterns and refinement strategies. Finally, our proposed CCS training brings the most substantial gains. For instance, on Qwen2.5-7B, CCS improves the average performance from 49.6% to 52.9%. These results highlight both the effectiveness of the CLM and the benefit of our contrastive alignment strategy. Detailed results are provided in Table 4 (Appendix).

### v4

# Retrieval is Not Enough: Enhancing RAG Reasoning through Test-Time Critique and Optimization

@@ Abstract
Retrieval-augmented generation (RAG) has become a widely adopted paradigm for enabling knowledge-grounded large language models (LLMs). However, standard RAG pipelines often fail to ensure that model reasoning remains consistent with the evidence retrieved, leading to factual inconsistencies or unsupported conclusions. In this work, we reinterpret RAG as Retrieval-Augmented Reasoning and identify a central but underexplored problem: Reasoning Misalignment—the divergence between an LLM’s internal reasoning trajectory and the evidential constraints provided by retrieval. To address this issue, we propose AlignRAG, a novel iterative framework grounded in Critique-Driven Alignment (CDA). We further introduce AlignRAG-auto, an autonomous variant that dynamically terminates refinement, removing the need to pre-specify the number of critique iterations. At the heart of AlignRAG lies a contrastive critique synthesis mechanism that generates retrieval-sensitive critiques while mitigating self-bias. This mechanism trains a dedicated retrieval-augmented Critic Language Model (CLM) using labeled critiques that distinguish between evidence-aligned and misaligned reasoning. Empirical evaluations show that our approach significantly improves reasoning fidelity. Our 8B-parameter CLM improves performance over the Self-Refine baseline by 12.1% on out-of-domain tasks and outperforms a standard 72B-parameter CLM by 2.2%. Furthermore, AlignRAG-auto achieves this state-of-the-art performance while dynamically determining the optimal number of refinement steps, enhancing efficiency and usability. AlignRAG remains compatible with existing RAG architectures as a plug-and-play module and demonstrates strong robustness under both informative and noisy retrieval scenarios. Overall, AlignRAG offers a principled solution for aligning model reasoning with retrieved evidence, substantially improving the factual reliability and robustness of RAG systems. Our source code is provided at Github.
@@ 1 Introduction
Large Language Models (LLMs) have significantly advanced natural language understanding and generation capabilities. Retrieval-Augmented Generation (RAG) [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] has emerged as a prominent paradigm for grounding LLM responses with external knowledge. However, RAG systems exhibit notable fragility, particularly when confronted with irrelevant or noisy retrieved evidence [11, 12]. Existing methods primarily rely on static, training-time optimizations, which are often insufficient to address the dynamic challenges of error propagation during inference [13, 14]. We identify a critical, yet understudied, failure mode in RAG: reasoning misalignment—a disconnect between the model’s reasoning process and the retrieved evidence. Prior work has focused predominantly on improving retrieval quality or generating more robust outputs, largely overlooking the explicit alignment of the reasoning steps with the provided evidence [15, 7, 16, 17, 18]. While reflective approaches like Self-RAG [19] attempt error detection, they often necessitate architectural modifications or task-specific fine-tuning, limiting their generalizability.
@@
To enhance the framework’s practicality, we also introduce AlignRAG-auto, a more autonomous variant that eliminates the need for manual tuning of iteration counts. By training the CLM to predict a special ‘[Good]‘ token upon generating a satisfactory response, AlignRAG-auto implements a dynamic stopping mechanism at inference time. This allows the system to terminate the refinement loop as soon as the reasoning is aligned with the evidence, saving substantial computational resources while maintaining high accuracy. This "hands-free" approach makes our framework more robust, efficient, and readily deployable across diverse tasks without domain-specific adjustments.
@@
In summary, this paper makes the following key contributions: (1) We reconceptualize RAG as Retrieval-Augmented Reasoning and identify Reasoning Misalignment as a fundamental, understudied failure mode. (2) We introduce critique learning for RAG, a novel pipeline for training CLMs to generate retrieval-augmented critiques while mitigating self-preference bias through a contrastive synthesis approach. (3) We propose AlignRAG, a test-time framework that utilizes CDA steps to iteratively optimize the RAG reasoning process. We also present AlignRAG-auto, an autonomous extension that dynamically determines the optimal number of refinement steps, enhancing efficiency and usability. (4) We provide extensive empirical validation demonstrating that AlignRAG achieves SOTA performance and significantly improves reasoning quality and robustness across various benchmarks and retrieval scenarios.
@@ 2 Reasoning Misalignment in RAG
Phase 3: Evidence-Integrated Synthesis. Misalignment here involves generating reasoning steps or conclusions $\text{step}_{i}\in\text{y}$ that are not logically supported by the relevant evidence $E_{i}\subseteq\mathcal{D}$ used for that step, violating the entailment $E_{i}\models\text{step}_{i}$, or creating inconsistencies when synthesizing information from multiple parts of $\mathcal{D}$ into a coherent justification [31, 32, 14].
@@ 3.2.1 Training Corpus Construction
To model the ambiguity and diversity inherent in real-world retrieval scenarios [33, 34, 35], we construct a structured training dataset $\mathcal{S}=\{(q_{i},a_{i},\mathcal{D}_{i},\mathbf{c}_{i})\}_{i=1}^{N}$, where each instance includes a query $q_{i}$, its gold answer $a_{i}$, a retrieved document set $\mathcal{D}_{i}$, and a context granularity vector $\mathbf{c}_{i}\in\{0,1\}^{3}$. The vector encodes three orthogonal axes of contextual variation:
@@ 3.3 Critique-Driven Alignment Systems
$\text{y}_{0}\xrightarrow[]{\text{CDA}}\text{y}_{1}\xrightarrow[]{\text{CDA}}\cdots\xrightarrow[]{\text{CDA}}\text{y}_{T},$ (7)
@@
3.4 AlignRAG-auto: Dynamic and Domain-General Alignment
We also develop AlignRAG-auto, an autonomous extension of CDA that eliminates the need for manual iteration tuning and demonstrates strong cross-domain generalization. Whereas the baseline CDA requires specifying the maximum number of refinement steps $T$, AlignRAG-auto leverages a lightweight control mechanism that dynamically determines when alignment has been achieved. This improves both efficiency and robustness at deployment.
3.4.1 Training for Dynamic Refinement
We modify the training process for the Critic Language Model (CLM) to support dynamic inference. For each “unexpected” response $y_{\text{unexp}}$, we first determine correctness relative to ground truth. Incorrect responses are labeled with a [Bad] token, while correct ones are labeled with [Good]. The CLM is then trained to generate the appropriate control token followed by a structured critique:
$p_{\theta}(\texttt{[Good/Bad]},\Delta y\mid q,\mathcal{D},y_{\text{unexp}}).$ (10)
This dual-target training objective equips the critic with both evaluative and corrective capacity. By conditioning critiques on explicit correctness judgments, the CLM learns to terminate refinement as soon as sufficient alignment is achieved.
3.4.2 Dynamic Inference-Time Stopping
At inference, AlignRAG-auto proceeds iteratively as in CDA but introduces a dynamic stopping rule. If the critic predicts [Good], the system halts refinement and accepts the current candidate:
$y_{t+1}=\begin{cases}y_{t},&\text{if CLM outputs {[Good]}}\\ \mathcal{M}_{\text{gen}}(y_{t}\oplus\Delta y_{t}),&\text{if CLM outputs {[Bad]}}.\end{cases}$ (11)
This adaptive strategy avoids unnecessary iterations, saving compute while maintaining high fidelity to retrieved evidence. Empirically, we observe that only a subset of responses require multiple refinements, yielding substantial runtime reductions without compromising accuracy (see Table 2).
@@ 4.1 Experiment Setup
Dataset. To train a strong critique generator, we construct a 10K dataset by sampling 2K instances from each of five benchmarks: PopQA [41], TriviaQA [42], NaturalQuestions [43], 2WikiMultihopQA [44], and ASQA [45]. Furthermore, we evaluate our method on the same five in-domain benchmarks, along with two out-of-distribution (OOD) tasks, i.e., HotpotQA [46] and SQuAD [47].
Baselines. In our experiments, we compare our method against a range of non-retrieval and retrieval-based baselines. For non-retrieval baselines, we include Chain-of-Thought (CoT) prompting [48, 49] applied to models without retrieval augmentation. For standard RAG, we report performance from Vanilla Reasoning [13, 50], which performs step-by-step answer generation based on the retrieved passages. To assess the benefits of intermediate supervision, we include training-time refinement baselines such as RetRobust[33] and InstructRAG[13]. Our main comparison is with these test-time refinement methods, as they share similar objectives. For test-time refinement, we evaluate Self-RAG [19] and Self-Refine, which iteratively revises outputs based on self-generated critique.
@@ 4.2 Main Result
Analysis of AlignRAG-auto.
The experimental results, detailed in Table 2, reveal a compelling comparison between the fixed-iteration and autonomous alignment strategies. Notably, AlignRAG-auto consistently achieves performance on par with, and in many cases slightly superior to, its AlignRAG-fixed counterpart across all datasets and model sizes. For instance, on the NQ and ASQA benchmarks, AlignRAG-auto demonstrates clear improvements, suggesting its dynamic termination mechanism is highly effective.
Figure 2: Performance Comparison: AlignRAG-fixed (1 iter.) vs. AlignRAG-auto
AlignRAG-fixed (1 iter.) AlignRAG-auto
Dataset 8B 7B 14B 8B 7B 14B
PopQA 66.5 68.4 68.4 67.6 68.1 68.3
TriviaQA 77.0 77.8 79.5 77.6 78.1 79.9
NQ 65.3 65.9 67.7 66.8 67.3 69.0
2WikiMultiHopQA 47.0 49.5 49.8 47.6 49.3 50.2
ASQA 47.1 48.9 48.6 48.8 49.6 49.8
This is a crucial finding: the autonomous variant does not sacrifice accuracy for efficiency. Instead, it demonstrates that the Critic Language Model is well-calibrated to dynamically determine the optimal number of refinement steps, leading to robust performance without the need for manual hyperparameter tuning. This validates AlignRAG-auto as a more practical and efficient framework for real-world deployment.
@@ 4.3 Analysis
Note: All subsequent experimental analyses are based on AlignRAG-fixed. For convenience, we’ll use AlignRAG instead.
@@
Figure 5: AlignRAG performance under Noisy Retrieval (a) and Informative Retrieval (b) scenarios.
@@
This dual strategy, as demonstrated in Figures 5(a) and 5(b), empowers robust reasoning even under these adversarial conditions. Consequently, AlignRAG (Figure 5(a)) significantly outperforms conventional methods like Vanilla RAG and Self-Refine, which falter in such scenarios, by effectively transforming imperfect external retrievals and inherent model understanding into actionable insights.
@@
We examine the effectiveness of using a high-capacity Strong LLM, i.e., Qwen2.5-72B-Instruct, in our Contrastive Critique Synthesis (CCS) framework. Specifically, we compare three CLM supervision strategies under three backbones: (1) Vanilla Critique Synthesis, where the CLM is trained with critiques directly generated by Qwen2.5-72B-Instruct, (2) directly using the 72B model as the CLM, and (3) our CCS-based CLM trained with critiques synthesized by the 72B Strong LLM. As shown in Figure 6, our method consistently outperforms vanilla critique synthesis, validating the benefit of introducing contrastive reasoning signals. Remarkably, our method even surpasses the directly supervised 72B CLM in all backbones, suggesting that contrastive critique training enhances generalization and reduces reliance on model scale. These results demonstrate the strength of our approach in leveraging powerful LLMs for scalable and efficient CLM training.
@@ 4.4 Ablation Study
We perform an ablation study on seven benchmarks to isolate the effect of the Critic Language Model (CLM) and our contrastive critique synthesis (CCS). Specifically, we compare four settings: (1) Vanilla RAG, which generates without critique; (2) RAG + Frozen CLM, where a pretrained LLaMA3.1-8B serves as an untrained critic; (3) RAG + Trained CLM, where the CLM is fine-tuned on critiques over weak LLM responses, but without contrastive signal; (4) RAG + Trained CLM (CCS), our full method with contrastive critique synthesis.
Figure 7 shows that even a frozen CLM consistently improves over Vanilla RAG, confirming the utility of auxiliary critique. Training the CLM with weak LLM critiques yields further gains by capturing recurring error patterns. The largest improvements come from CCS, which explicitly contrasts expert and weak responses. For example, on Qwen2.5-7B, CCS raises average accuracy from 49.6% to 52.9%. Complete results are presented in Table 4 (Appendix).
5 Related Work
RAG [53, 54, 55, 56, 57, 58, 59] enhances LLMs by grounding generation in external knowledge. While significant efforts have focused on improving retrieval accuracy [60, 61, 62, 63, 64] or training better generators [65, 66, 67, 68], a persistent challenge is the presence of noisy or irrelevant retrieved content. To mitigate this, recent methods filter or denoise context, predominantly via training-time optimizations [15, 69, 70, 71, 72], such as InstructRAG’s [13] self-supervised denoising. However, these approaches are limited by their static nature and fail to address dynamic error propagation during inference [14].